OpenStudy (anonymous):

Find the intervals of increase and decrease of 1/x^2-9 b) find the coordinates of the local max and min Please help, I got x=0 and x=+-3 but im not sure what the intervals would be for the increase and decrease

OpenStudy (michele_laino):

the first derivative of your function is: \[\Large f'\left( x \right) = \frac{{ - 2x}}{{{{\left( {{x^2} - 9} \right)}^2}}}\]

OpenStudy (michele_laino):

am I right?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! Now we have to establish where that first derivative is positive

OpenStudy (michele_laino):

So we have to solve this inequality: \[\Large \frac{{ - 2x}}{{{{\left( {{x^2} - 9} \right)}^2}}} > 0\]

OpenStudy (michele_laino):

what are the solutions of that inequality?

OpenStudy (anonymous):

would you set the denominator to = 0? i got +-3

OpenStudy (michele_laino):

no, since the denominator is always positive, then that fraction will be positive, if and only if also the numerator will be positive, namely: \[\Large - 2x > 0\]

OpenStudy (michele_laino):

the denominator: \[\Large {{{\left( {{x^2} - 9} \right)}^2}}\] is a square, which is always positive, and it is equal to zero at x=3 and x=-3 That means our first derivative doesn't exist at x=3 and x=-3. Also your function doesn't exist at x=3 and x=-3 for the same reason

OpenStudy (anonymous):

Oh okay that makes sense

OpenStudy (michele_laino):

ok! Now what are the solutions of: \[\Large - 2x > 0\]

OpenStudy (anonymous):

it would be 0

OpenStudy (michele_laino):

if I multiply, both sides of that inequality by -1, I get: \[\Large 2x < 0\] am I right?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! Now please, divide both sides of the last inequality by 2, what do you get?

OpenStudy (anonymous):

x<0 (0/2)

OpenStudy (michele_laino):

perfect!

OpenStudy (michele_laino):

that means our function is an incresing function when x<0 and a decreasing function when x>0: 

OpenStudy (michele_laino):

increasing*

OpenStudy (michele_laino):

so what can you conclude?

OpenStudy (anonymous):

Okay, would my intervals be then increasing from (-infinity,0) and decreasingfrom (0,-infinity)?

OpenStudy (michele_laino):

that's right! Please keep in mind that at subsequent points: x=-3 and x=3, our function is not defined

OpenStudy (michele_laino):

your last statement means that x=0 is a point of local maximum for f(x), am I right?

OpenStudy (michele_laino):

decreasing in (0, +infinity) and increasing in (-infinity, 0)

OpenStudy (anonymous):

thats what i have, but Im not sure if its right, would the -3 and 3 have anything to do with the local max and min?

OpenStudy (michele_laino):

no, they are not. Those points don't belong to the domain of our function